OpenStudy (anonymous):

Find the area of a section of a circle which has 8" radius if the sector has an arc that measures 45°. _____ times pi sq. in.

OpenStudy (kropot72):

Area of sector is \[1/2 \times r ^{2}\times \theta\]

OpenStudy (anonymous):

if area = pi r2, then 16 * 16* 256, 804, now lets cut this circle into eight pieces, like this, they are cut into 8 pieces with 2 45, circumference lines, in a manner we can say that 804/8 will give you your answer, which is 100.5

OpenStudy (anonymous):



OpenStudy (anonymous):

That is what you where asking for, am I correct?

OpenStudy (kropot72):

Area of sector is:\[((\pi \times r ^{2})\div360)\times45=25.13\square inches\]

OpenStudy (anonymous):

Thanks everyone [:

OpenStudy (anonymous):

I hope I was right, correct me if im not, and as always, my plesure :D

OpenStudy (kropot72):

The correct form of the answer is : \[8\times\pi) squareinches\]